OpenStudy (usukidoll):

separable equation 1/x dx + dy = 0

hartnn (hartnn):

easy for you

OpenStudy (usukidoll):

\[\frac{1}{x}+dy = 0\]

OpenStudy (usukidoll):

\[dy = - \frac{1}{x} dx\]

hartnn (hartnn):

yeah, now integrate

OpenStudy (usukidoll):

\[y = - \ln x + c\]

hartnn (hartnn):

correct! :)

OpenStudy (usukidoll):

\[y + \ln x = c\]

OpenStudy (usukidoll):

I need to refresh my memory thanks to pure math. I'm trying to get back to the computation parts XD.

OpenStudy (anonymous):